OpenStudy (anonymous):

Solve the equation. 2^(x) +3.2^(x+2) + 2^(x+3) = 84

OpenStudy (hunus):

\[2^x +3*2^{x+2} + 2^{x+3} = 84\] \[2^x(1+3*2^2+2^3)=84\] \[2^x(21)=84\] \[2^x=4\] \[x=2\]

OpenStudy (anonymous):

x=2

OpenStudy (saifoo.khan):

are u sure that's correct?

OpenStudy (saifoo.khan):

coz it's 3.2

OpenStudy (hunus):

I read it as 3*2 o.o

OpenStudy (anonymous):

it is 3*2

OpenStudy (saifoo.khan):

then it's correct.